OpenStudy (anonymous):

What is 2/3 divided by 5/6?

OpenStudy (anonymous):

Have you ever heard of the rule keep it switch it flip it?

OpenStudy (mathstudent55):

To divide a fraction by a fraction: Copy the first fraction as it is. Change the division to a multiplication. Flip the second fraction. Then multiply the fractions. Example: \(\dfrac{3}{5} \div \dfrac{7}{8} = \dfrac{3}{5} \color{red}\times \color{green}{\dfrac{8}{7}} \) \(= \dfrac{3 \times 8}{5 \times 7} = \dfrac{24}{35} \)

OpenStudy (mathstudent55):

@mycars12 I never heard of the rule, but I'll remember it now.

OpenStudy (anonymous):

ohh i learned it in algebra 1 lol, we thought my teacher was crazy but it stuck in our heads

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }\div \frac{ 5 }{ 6 }\] Multiply by the flipped fraction, simplify if possible. \[\frac{ 2 }{ 3 }\times \frac{ 6 }{ 5 }=\frac{ 18 }{ 10 }\]

OpenStudy (mathstudent55):

@jcpd910 Wait. 2 * 6 = 18??? 3 * 5 = 10 ???

OpenStudy (anonymous):

You cross multiply.

OpenStudy (anonymous):

Wait, no you don't! xD it's just 12 over 15